You have taken a job as an entry-level analyst, and your boss has asked you to find the expected value of Pandar Corp's stock. As you were doing your research, you found out that Pandar Corp just paid a dividened (Do) of $3.75. The firm has experienced consistent growth of 5% for the last couple of years, and you belive that the firm will continue to grow at the same rate in the future.

1. If investors require a return of 10% on Pander Corp's stock, what is the expected value of the company's stock?

2. What would be the change in the expected value of Pander Corp's stock if investors required a return of 12% on the company's stock?
Answer #1


D1 = Do * (1+g) = 3.75 * (1+0.05) = 3.9375


expected value of stock = D1/(r-g) = 3.9375/(10%-5%) = 78.75


2) if r = 12%


expected value of stock = 3.9375/(12% - 5%) = 56.25



chenge in expected value of stock = 78.75 - 56.25/78.75 * 100 = 28.57%